# TabMGP: Martingale Posterior with TabPFN

### v1

@@ Appendix F Diagnostics
where $p\coloneq\dim(\theta)$ and $\theta(F_{n})$ is the empirical risk minimiser (or maximum likelihood estimate of $z_{1:n}$ in our experiment). In practice, we simulate many realisations of $F_{N}$ to approximate this expectation. This expectation is shown in Figure 1 and corresponds to the blue trajectories in Figure 5. In addition to the expected value, we also present the $L_{1}$-norm for each realisation of $F_{N}$ in Figure 5. While the paths corresponding to individual realisations of $F_{N}$ exhibit variability, they generally stabilise as $N$ increases, providing further empirical evidence that $\theta(F_{N})$ converges $\mathbb{P}$}-a.s..
@@
Figure 5: $L_{1}$-norm between $\theta(F_{n})$ and $\theta(F_{N})$ as $N$ increases. Each of the 30 plots corresponds to a realisation $z_{1:n}$ of a particular setup. Each of the grey lines corresponds to a realisation of $F_{N}$, and the blue lines are their average.
@@ Appendix G Marginal density and credible intervals of the posteriors
Figure 6: Marginal density plots of linear regression with the synthetic Gaussian i.i.d. data. For each posterior density, the 95% marginal credible interval is shown as a horizontal bar, and the posterior mean as dots. The vertical dotted red and blue lines correspond to $\theta(F_{0})$ and $\theta(F_{n})$, respectively.
Figure 7: Marginal density plots of linear regression with the synthetic Bernoulli data with a logistic link. For each posterior density, the 95% marginal credible interval is shown as a horizontal bar, and the posterior mean as dots. The vertical dotted red and blue lines correspond to $\theta(F_{0})$ and $\theta(F_{n})$, respectively.
Figure 8: Marginal density plots for the ‘abalone’ dataset. For each posterior density, the 95% marginal credible interval is shown as a horizontal bar, and the posterior mean as dots. The vertical dotted red and blue lines correspond to $\theta(F_{0})$ and $\theta(F_{n})$, respectively.
Figure 9: Marginal density plots for the ‘telescope’ dataset. For each posterior density, the 95% marginal credible interval is shown as a horizontal bar, and the posterior mean as dots. The vertical dotted red and blue lines correspond to $\theta(F_{0})$ and $\theta(F_{n})$, respectively.
Figure 10: Marginal density plots for the ‘yeast’ dataset. For each posterior density, the 95% marginal credible interval is shown as a horizontal bar, and the posterior mean as dots. The vertical dotted red and blue lines correspond to $\theta(F_{0})$ and $\theta(F_{n})$, respectively.
@@ Appendix H Coverage rate of the marginal credible interval
In addition to the joint credible sets discussed in the main text, we evaluate the performance of the $95\%$ marginal credible intervals (13) for each parameter. Figure 11 shows boxplots of the marginal coverage rates for each method across all parameters in our experiments. While most methods’ coverage is centered near the nominal 0.95 level, BB exhibits substantial under-coverage.
To jointly assess both coverage (calibration) and interval width (sharpness), we use the Winkler score (Winkler, 1972). For a $(1-\alpha)$ interval $[l,u]$ and a true value $\theta_{0}$, the score is defined as:
$W_{\alpha}(l,u,\theta_{0})=(u-l)+\frac{2}{\alpha}(l-\theta_{0}){\bm{1}}\{\theta_{0}<l\}+\frac{2}{\alpha}(\theta_{0}-u){\bm{1}}\{\theta_{0}>u\}.$
The score penalises wide intervals and adds a larger penalty if the interval fails to cover the true value. A lower score indicates better performance.
The subsequent tables (Table 5 to Table 12) shows the empirical coverage and the median Winkler score (in parentheses) for each parameter, computed over 100 data realisations. The method with the lowest median Winkler score for each parameter is highlighted in bold. Summarising across all 30 setups and 293 individual parameters, TabMGP (TabPFN) achieves the best score 155 times, followed by Bayesian bootstrap (125), Copula (11), and the classical Bayesian posterior (2). TabPFN performs particularly well in logistic regression settings, while the Bayesian bootstrap is often competitive in linear regression, though its intervals frequently under-cover.
@@
$z_{1:n}$ Feature Name TabMGP $B$ Copula Bayes
$N(0,1)$ intercept 0.97 (0.73) 0.83 (0.34) 1.00 (0.63) 1.00 (1.24)
x1 1.00 (0.64) 0.77 (0.38) 1.00 (0.67) 1.00 (1.34)
x2 0.94 (0.68) 0.75 (0.37) 0.99 (0.65) 1.00 (1.26)
x3 0.99 (0.65) 0.76 (0.37) 0.99 (0.65) 1.00 (1.28)
x4 0.99 (0.64) 0.81 (0.36) 0.99 (0.65) 1.00 (1.28)
x5 0.99 (0.66) 0.76 (0.38) 1.00 (0.67) 1.00 (1.29)
x6 0.95 (0.70) 0.72 (0.38) 0.97 (0.67) 1.00 (1.30)
x7 0.96 (0.69) 0.67 (0.39) 0.98 (0.66) 1.00 (1.23)
x8 0.94 (0.73) 0.78 (0.36) 0.96 (0.67) 1.00 (1.27)
x9 0.97 (0.71) 0.79 (0.36) 0.98 (0.67) 1.00 (1.26)
x10 0.97 (0.80) 0.79 (0.38) 0.98 (0.63) 1.00 (1.28)
$t_{5}$ intercept 0.97 (0.80) 0.79 (0.40) 1.00 (0.65) 1.00 (1.24)
x1 1.00 (0.67) 0.78 (0.43) 0.99 (0.67) 1.00 (1.33)
x3 0.99 (0.66) 0.76 (0.43) 1.00 (0.65) 1.00 (1.27)
x4 1.00 (0.66) 0.77 (0.40) 0.97 (0.66) 1.00 (1.28)
x5 0.99 (0.67) 0.75 (0.42) 0.99 (0.68) 1.00 (1.28)
x6 0.98 (0.75) 0.76 (0.46) 0.97 (0.67) 1.00 (1.29)
x7 0.99 (0.70) 0.78 (0.42) 1.00 (0.66) 1.00 (1.23)
x8 0.94 (0.75) 0.68 (0.43) 0.97 (0.66) 1.00 (1.26)
x9 0.97 (0.75) 0.81 (0.41) 0.98 (0.67) 1.00 (1.26)
x10 0.92 (0.86) 0.76 (0.44) 0.88 (0.69) 1.00 (1.28)
$t_{4}$ intercept 0.97 (0.81) 0.81 (0.41) 1.00 (0.64) 1.00 (1.23)
x1 1.00 (0.68) 0.79 (0.46) 0.99 (0.67) 1.00 (1.33)
x2 0.94 (0.71) 0.76 (0.47) 0.96 (0.68) 1.00 (1.25)
x3 0.99 (0.67) 0.78 (0.45) 1.00 (0.65) 1.00 (1.27)
x4 1.00 (0.68) 0.78 (0.43) 0.98 (0.66) 1.00 (1.29)
x5 0.99 (0.67) 0.76 (0.45) 0.99 (0.67) 1.00 (1.28)
x6 0.99 (0.75) 0.75 (0.49) 0.97 (0.67) 1.00 (1.30)
x7 0.98 (0.72) 0.78 (0.43) 0.98 (0.66) 1.00 (1.24)
x8 0.94 (0.76) 0.70 (0.46) 0.97 (0.65) 1.00 (1.26)
x9 0.97 (0.74) 0.81 (0.43) 0.98 (0.67) 1.00 (1.27)
x10 0.93 (0.87) 0.75 (0.48) 0.87 (0.69) 1.00 (1.28)
$t_{3}$ intercept 0.97 (0.81) 0.82 (0.43) 0.96 (0.63) 0.99 (1.24)
x1 1.00 (0.68) 0.79 (0.48) 0.99 (0.66) 1.00 (1.34)
x2 0.95 (0.70) 0.76 (0.49) 0.96 (0.68) 1.00 (1.25)
x3 0.99 (0.68) 0.78 (0.49) 0.98 (0.65) 0.99 (1.27)
x4 1.00 (0.68) 0.78 (0.51) 0.98 (0.66) 0.99 (1.29)
x5 0.99 (0.67) 0.73 (0.51) 0.98 (0.66) 1.00 (1.28)
x6 0.98 (0.73) 0.73 (0.54) 0.94 (0.66) 0.98 (1.33)
x7 0.98 (0.70) 0.77 (0.45) 0.98 (0.65) 0.98 (1.25)
x8 0.92 (0.73) 0.71 (0.47) 0.93 (0.65) 0.99 (1.27)
x9 0.98 (0.74) 0.79 (0.45) 0.97 (0.66) 1.00 (1.26)
x10 0.93 (0.86) 0.74 (0.52) 0.85 (0.68) 1.00 (1.28)
@@
Table 5: Coverage and Winkler score (median over 100 repetitions, shown in parantheses) of the marginal credible intervals for linear regression with synthetic i.i.d. data. The method with the lowest Winkler score is shown in bold. The target coverage is 0.95.
$z_{1:n}$ Feature Name TabMGP $B$ Copula Bayes
$s_{1}$ intercept 0.99 (0.66) 0.79 (0.21) 1.00 (0.63) 1.00 (1.24)
x1 1.00 (0.55) 0.73 (0.24) 1.00 (0.67) 1.00 (1.34)
x2 0.97 (0.61) 0.73 (0.22) 1.00 (0.65) 1.00 (1.25)
x3 1.00 (0.59) 0.80 (0.22) 1.00 (0.66) 1.00 (1.27)
x4 1.00 (0.57) 0.77 (0.23) 1.00 (0.66) 1.00 (1.28)
x5 1.00 (0.59) 0.77 (0.23) 1.00 (0.66) 1.00 (1.29)
x6 0.98 (0.64) 0.67 (0.25) 0.99 (0.65) 1.00 (1.31)
x7 0.99 (0.59) 0.65 (0.25) 1.00 (0.64) 1.00 (1.23)
x8 0.97 (0.64) 0.72 (0.23) 0.98 (0.67) 1.00 (1.26)
x9 0.99 (0.62) 0.78 (0.23) 1.00 (0.66) 1.00 (1.27)
x10 0.99 (0.71) 0.78 (0.21) 0.98 (0.64) 1.00 (1.29)
$s_{2}$ intercept 1.00 (0.65) 0.82 (0.15) 1.00 (0.63) 1.00 (1.24)
x1 0.99 (0.55) 0.62 (0.24) 1.00 (0.66) 1.00 (1.34)
x2 0.99 (0.58) 0.78 (0.19) 1.00 (0.65) 1.00 (1.25)
x3 1.00 (0.55) 0.72 (0.19) 1.00 (0.65) 1.00 (1.27)
x4 1.00 (0.55) 0.76 (0.19) 1.00 (0.67) 1.00 (1.28)
x5 0.99 (0.58) 0.76 (0.20) 1.00 (0.67) 1.00 (1.28)
x6 1.00 (0.61) 0.64 (0.22) 0.99 (0.65) 1.00 (1.31)
x7 0.99 (0.57) 0.75 (0.18) 1.00 (0.63) 1.00 (1.24)
x8 0.98 (0.61) 0.66 (0.20) 1.00 (0.66) 1.00 (1.27)
x9 1.00 (0.60) 0.77 (0.18) 1.00 (0.65) 1.00 (1.26)
x10 0.99 (0.70) 0.75 (0.18) 0.96 (0.64) 1.00 (1.28)
$s_{3}$ intercept 1.00 (0.63) 0.82 (0.14) 1.00 (0.64) 1.00 (1.23)
x1 0.99 (0.55) 0.62 (0.24) 1.00 (0.66) 1.00 (1.34)
x2 0.98 (0.59) 0.78 (0.17) 1.00 (0.65) 1.00 (1.25)
x3 0.99 (0.56) 0.68 (0.18) 1.00 (0.66) 1.00 (1.27)
x4 1.00 (0.57) 0.69 (0.21) 1.00 (0.67) 1.00 (1.29)
x5 1.00 (0.58) 0.75 (0.18) 1.00 (0.67) 1.00 (1.29)
x6 1.00 (0.61) 0.62 (0.21) 1.00 (0.65) 1.00 (1.30)
x7 1.00 (0.56) 0.76 (0.17) 1.00 (0.63) 1.00 (1.23)
x8 0.98 (0.61) 0.67 (0.19) 1.00 (0.66) 1.00 (1.27)
x9 1.00 (0.61) 0.74 (0.17) 1.00 (0.66) 1.00 (1.27)
x10 0.99 (0.71) 0.74 (0.18) 0.97 (0.66) 1.00 (1.28)
@@
Table 6: Coverage and Winkler score (median over 100 repetitions, shown in parantheses) of the marginal credible intervals for linear regression with synthetic, dependent-error data. The method with the lowest Winkler score is shown in bold. The target coverage is 0.95.
$z_{1:n}$ Feature Name TabMGP $B$ Copula Bayes
Logistic intercept_1.0 0.96 (1.11) 0.88 (1.40) 0.98 (1.67) 0.92 (1.46)
x1_1.0 0.97 (0.97) 0.86 (1.45) 0.99 (1.70) 0.87 (1.48)
x2_1.0 0.85 (1.37) 0.76 (1.59) 0.98 (1.84) 0.79 (1.58)
x3_1.0 0.95 (1.11) 0.81 (1.48) 1.00 (1.69) 0.85 (1.53)
x4_1.0 0.98 (0.99) 0.85 (1.44) 0.99 (1.71) 0.85 (1.49)
x5_1.0 0.94 (1.05) 0.86 (1.47) 0.99 (1.74) 0.88 (1.50)
x6_1.0 0.90 (1.41) 0.77 (1.62) 1.00 (1.88) 0.80 (1.62)
x7_1.0 0.89 (1.31) 0.79 (1.61) 1.00 (1.85) 0.82 (1.58)
x8_1.0 0.85 (1.59) 0.75 (1.84) 0.99 (2.01) 0.81 (1.75)
x9_1.0 0.89 (1.58) 0.73 (1.70) 0.99 (1.94) 0.76 (1.70)
x10_1.0 0.86 (1.88) 0.66 (2.12) 0.99 (2.26) 0.71 (2.05)
GMM(0)(0) intercept_1.0 0.90 (1.54) 0.73 (2.09) 0.99 (2.18) 0.82 (2.06)
x1_1.0 0.96 (1.21) 0.82 (1.90) 1.00 (1.98) 0.89 (1.83)
x2_1.0 0.86 (1.77) 0.70 (2.21) 0.97 (2.17) 0.77 (2.04)
x3_1.0 0.91 (1.44) 0.79 (1.92) 0.99 (1.96) 0.84 (1.87)
x4_1.0 0.97 (1.26) 0.79 (1.85) 0.99 (2.00) 0.79 (1.89)
x5_1.0 0.94 (1.26) 0.84 (1.81) 0.97 (1.90) 0.87 (1.86)
x6_1.0 0.91 (1.96) 0.68 (2.51) 1.00 (2.37) 0.70 (2.23)
x7_1.0 0.93 (1.73) 0.74 (2.20) 0.99 (2.11) 0.78 (2.05)
x8_1.0 0.90 (2.30) 0.62 (2.97) 0.99 (2.51) 0.66 (2.52)
x9_1.0 0.95 (1.94) 0.64 (2.79) 1.00 (2.33) 0.69 (2.38)
x10_1.0 0.91 (2.74) 0.52 (3.92) 1.00 (3.00) 0.57 (3.18)
GMM(1)(-1) intercept_1.0 0.96 (1.10) 0.89 (1.40) 0.98 (1.67) 0.92 (1.46)
x1_1.0 0.96 (0.96) 0.82 (1.47) 1.00 (1.68) 0.86 (1.52)
x2_1.0 0.85 (1.33) 0.77 (1.59) 0.99 (1.81) 0.84 (1.60)
x3_1.0 0.91 (1.13) 0.83 (1.55) 0.99 (1.68) 0.89 (1.50)
x4_1.0 0.96 (0.99) 0.82 (1.46) 0.99 (1.69) 0.87 (1.48)
x5_1.0 0.93 (1.05) 0.88 (1.46) 1.00 (1.71) 0.90 (1.50)
x6_1.0 0.88 (1.43) 0.75 (1.64) 1.00 (1.86) 0.81 (1.63)
x7_1.0 0.85 (1.30) 0.78 (1.59) 1.00 (1.82) 0.83 (1.54)
x8_1.0 0.86 (1.66) 0.68 (1.87) 0.99 (2.05) 0.73 (1.79)
x9_1.0 0.94 (1.48) 0.73 (1.74) 1.00 (1.92) 0.73 (1.71)
x10_1.0 0.90 (1.81) 0.66 (2.06) 1.00 (2.25) 0.73 (2.02)
GMM(2)(-2) intercept_1.0 0.92 (1.06) 0.84 (1.37) 0.98 (1.68) 0.88 (1.38)
x1_1.0 0.97 (0.81) 0.89 (1.24) 0.97 (1.57) 0.93 (1.32)
x2_1.0 0.87 (1.00) 0.78 (1.30) 0.99 (1.60) 0.83 (1.37)
x3_1.0 0.94 (0.89) 0.88 (1.27) 0.97 (1.58) 0.89 (1.33)
x4_1.0 0.99 (0.84) 0.91 (1.24) 1.00 (1.54) 0.92 (1.32)
x5_1.0 0.90 (0.91) 0.82 (1.32) 1.00 (1.53) 0.86 (1.32)
x6_1.0 0.91 (1.15) 0.84 (1.34) 1.00 (1.67) 0.84 (1.40)
x7_1.0 0.83 (1.04) 0.84 (1.37) 1.00 (1.61) 0.87 (1.34)
x8_1.0 0.81 (1.33) 0.78 (1.43) 1.00 (1.74) 0.82 (1.44)
x9_1.0 0.83 (1.20) 0.79 (1.35) 1.00 (1.65) 0.79 (1.41)
x10_1.0 0.85 (1.53) 0.73 (1.57) 1.00 (1.82) 0.80 (1.61)

### v2

@@ Appendix F Diagnostics
where $p\coloneq\dim(\theta)$ and $\theta(F_{n})$ is the empirical risk minimiser (or maximum likelihood estimate of $z_{1:n}$ in our experiment). In practice, we simulate many realisations of $F_{N}$ to approximate this expectation. This expectation is shown in Figure 3 and corresponds to the black solid trajectories in Figure 6. In addition to the expected value, we also present the $L_{1}$-norm for each realisation of $F_{N}$ in Figure 6. While the paths corresponding to individual realisations of $F_{N}$ exhibit variability, they generally stabilise as $N$ increases, providing further empirical evidence that $\theta(F_{N})$ converges $\mathbb{P}$}-a.s..
@@
Figure 6: $L_{1}$-norm between $\theta(F_{n})$ and $\theta(F_{N})$ as $N$ increases. Each of the 30 plots corresponds to a realisation $z_{1:n}$ of a particular setup. Each of the grey lines corresponds to a realisation of $F_{N}$, and the black solid lines are their average.
@@ Appendix G Marginal density and credible intervals of the posteriors
Figure 7: Marginal density plots of linear regression with the synthetic Gaussian i.i.d. data. For each posterior density, the 95% marginal credible interval is shown as a horizontal bar, and the posterior mean is marked. The solid and dashed black vertical lines correspond to $\theta(F^{\star})$ and $\theta(F_{n})$ respectively.
Figure 8: Marginal density plots of linear regression with the synthetic Bernoulli data with a logistic link. For each posterior density, the 95% marginal credible interval is shown as a horizontal bar, and the posterior mean is marked. The solid and dashed black vertical lines correspond to $\theta(F^{\star})$ and $\theta(F_{n})$ respectively.
Figure 9: Marginal density plots for the ‘abalone’ dataset. For each posterior density, the 95% marginal credible interval is shown as a horizontal bar, and the posterior mean is marked. The solid and dashed black vertical lines correspond to $\theta(F^{\star})$ and $\theta(F_{n})$ respectively.
Figure 10: Marginal density plots for the ‘telescope’ dataset. For each posterior density, the 95% marginal credible interval is shown as a horizontal bar, and the posterior mean is marked. The solid and dashed black vertical lines correspond to $\theta(F^{\star})$ and $\theta(F_{n})$ respectively.
Figure 11: Marginal density plots for the ‘yeast’ dataset. For each posterior density, the 95% marginal credible interval is shown as a horizontal bar, and the posterior mean is marked. The solid and dashed black vertical lines correspond to $\theta(F^{\star})$ and $\theta(F_{n})$ respectively.
@@ Appendix H Coverage rate of the marginal credible interval
In addition to the joint credible sets discussed in the main text, we evaluate the performance of the $95\%$ marginal credible intervals (13) for each parameter. Figure 12 shows boxplots of the marginal coverage rates for each method across all parameters in our experiments. While most methods’ coverage is centred near the nominal 0.95 level, BB exhibits substantial under-coverage. Bayes also tends to undercover in logistic regression.
@@
The subsequent tables (Table 5 to Table 12) shows the empirical coverage and the median width for each parameter, computed over 100 data realisations. The method achieving at least the target coverage with the smallest size is highlighted in bold. If none of the methods achieve the target coverage, the method with the highest coverage is highlighted in bold instead. Summarising across all 30 setups and 293 individual parameters, TabMGP is bolded 152 times, followed by Copula (100), the classical Bayesian posterior (37) and Bayesian bootstrap (4). TabMGP performs particularly well in multinomial logistic regression relative to the competing methods.
@@
Table 5: Coverage (rate) and median width (size) over 100 repetitions of marginal credible intervals for linear regression with synthetic i.i.d. data. The method achieving at least the target coverage with the smallest size is shown in bold. If none of the methods achieve the target coverage, the method with the highest coverage is shown in bold. The target coverage is 0.95.
TabMGP $B$ Copula Bayes
Setup Feature Name Rate Size Rate Size Rate Size Rate Size
intercept 0.97 0.73 0.83 0.30 1.00 0.63 1.00 0.87
x1 1.00 0.64 0.77 0.31 1.00 0.67 1.00 0.94
x10 0.97 0.78 0.79 0.28 0.98 0.62 1.00 0.90
x2 0.94 0.65 0.75 0.30 0.99 0.65 1.00 0.88
x3 0.99 0.65 0.76 0.32 0.99 0.65 1.00 0.90
x4 0.99 0.63 0.81 0.29 0.99 0.64 1.00 0.90
x5 0.99 0.65 0.76 0.31 1.00 0.67 1.00 0.91
x6 0.95 0.68 0.72 0.31 0.97 0.65 1.00 0.92
x7 0.96 0.68 0.67 0.30 0.98 0.65 1.00 0.86
x8 0.94 0.69 0.78 0.30 0.96 0.66 1.00 0.90
$N(0,1)$ x9 0.97 0.70 0.79 0.33 0.98 0.67 1.00 0.90
intercept 0.97 0.78 0.79 0.34 1.00 0.65 1.00 0.87
x1 1.00 0.67 0.78 0.36 0.99 0.67 1.00 0.95
x10 0.92 0.84 0.76 0.37 0.88 0.63 1.00 0.90
x2 0.96 0.70 0.76 0.37 0.96 0.67 0.99 0.89
x3 0.99 0.66 0.76 0.36 1.00 0.65 1.00 0.91
x4 1.00 0.66 0.77 0.34 0.97 0.65 1.00 0.91
x5 0.99 0.66 0.75 0.35 0.99 0.68 1.00 0.91
x6 0.98 0.74 0.76 0.36 0.97 0.66 1.00 0.92
x7 0.99 0.69 0.78 0.36 1.00 0.66 1.00 0.87
x8 0.94 0.73 0.68 0.34 0.97 0.64 1.00 0.89
$t_{5}$ x9 0.97 0.74 0.81 0.34 0.98 0.66 1.00 0.90
intercept 0.97 0.80 0.81 0.35 1.00 0.64 1.00 0.87
x1 1.00 0.68 0.79 0.38 0.99 0.67 1.00 0.95
x10 0.93 0.83 0.75 0.40 0.87 0.63 0.99 0.90
x2 0.94 0.68 0.76 0.39 0.96 0.66 0.99 0.88
x3 0.99 0.67 0.78 0.38 1.00 0.65 1.00 0.90
x4 1.00 0.68 0.78 0.36 0.98 0.65 1.00 0.92
x5 0.99 0.67 0.76 0.36 0.99 0.66 1.00 0.90
x6 0.99 0.75 0.75 0.38 0.97 0.67 1.00 0.91
x7 0.98 0.71 0.78 0.38 0.98 0.65 0.99 0.87
x8 0.94 0.71 0.70 0.34 0.97 0.64 0.99 0.89
$t_{4}$ x9 0.97 0.72 0.81 0.36 0.98 0.67 1.00 0.89
intercept 0.97 0.80 0.82 0.38 0.96 0.62 0.99 0.87
x1 1.00 0.68 0.79 0.40 0.99 0.66 0.99 0.95
x10 0.93 0.83 0.74 0.41 0.85 0.62 0.98 0.91
x2 0.95 0.67 0.76 0.42 0.96 0.66 0.97 0.89
x3 0.99 0.67 0.78 0.40 0.98 0.65 0.98 0.90
x4 1.00 0.68 0.78 0.38 0.98 0.65 0.98 0.90
x5 0.99 0.66 0.73 0.39 0.98 0.66 1.00 0.91
x6 0.98 0.73 0.73 0.41 0.94 0.65 0.96 0.92
x7 0.98 0.69 0.77 0.40 0.98 0.64 0.98 0.87
x8 0.92 0.71 0.71 0.37 0.93 0.62 0.97 0.90
$t_{3}$ x9 0.98 0.73 0.79 0.38 0.97 0.66 0.97 0.90
@@
Table 6: Coverage (rate) and median width (size) over 100 repetitions of marginal credible intervals for linear regression with synthetic, dependent-error data. The method achieving at least the target coverage with the smallest size is shown in bold. If none of the methods achieve the target coverage, the method with the highest coverage is shown in bold. The target coverage is 0.95.
TabMGP $B$ Copula Bayes
Setup Feature Name Rate Size Rate Size Rate Size Rate Size
intercept 0.99 0.66 0.79 0.18 1.00 0.63 1.00 0.86
x1 1.00 0.55 0.73 0.19 1.00 0.67 1.00 0.94
x10 0.99 0.71 0.78 0.19 0.98 0.64 1.00 0.91
x2 0.97 0.61 0.73 0.18 1.00 0.65 1.00 0.89
x3 1.00 0.59 0.80 0.20 1.00 0.66 1.00 0.90
x4 1.00 0.57 0.77 0.19 1.00 0.66 1.00 0.91
x5 1.00 0.59 0.77 0.20 1.00 0.66 1.00 0.91
x6 0.98 0.64 0.67 0.19 0.99 0.65 1.00 0.92
x7 0.99 0.59 0.65 0.19 1.00 0.64 1.00 0.87
x8 0.97 0.63 0.72 0.19 0.98 0.66 1.00 0.89
$s_{1}$ x9 0.99 0.62 0.78 0.20 1.00 0.66 1.00 0.89
intercept 1.00 0.65 0.82 0.14 1.00 0.63 1.00 0.86
x1 0.99 0.54 0.62 0.16 1.00 0.66 1.00 0.95
x10 0.99 0.70 0.75 0.15 0.96 0.64 1.00 0.91
x2 0.99 0.58 0.78 0.14 1.00 0.65 1.00 0.89
x3 1.00 0.55 0.72 0.15 1.00 0.65 1.00 0.90
x4 1.00 0.55 0.76 0.15 1.00 0.67 1.00 0.91
x5 0.99 0.58 0.76 0.16 1.00 0.67 1.00 0.91
x6 1.00 0.61 0.64 0.15 0.99 0.65 1.00 0.92
x7 0.99 0.57 0.75 0.15 1.00 0.63 1.00 0.87
x8 0.98 0.61 0.66 0.15 1.00 0.66 1.00 0.89
$s_{2}$ x9 1.00 0.60 0.77 0.16 1.00 0.65 1.00 0.90
intercept 1.00 0.63 0.82 0.13 1.00 0.64 1.00 0.87
x1 0.99 0.55 0.62 0.15 1.00 0.66 1.00 0.95
x10 0.99 0.70 0.74 0.14 0.97 0.64 1.00 0.90
x2 0.98 0.58 0.78 0.13 1.00 0.65 1.00 0.89
x3 0.99 0.56 0.68 0.15 1.00 0.66 1.00 0.91
x4 1.00 0.57 0.69 0.14 1.00 0.67 1.00 0.91
x5 1.00 0.58 0.75 0.15 1.00 0.67 1.00 0.90
x6 1.00 0.61 0.62 0.15 1.00 0.65 1.00 0.92
x7 1.00 0.56 0.76 0.14 1.00 0.63 1.00 0.87
x8 0.98 0.60 0.67 0.14 1.00 0.66 1.00 0.90
$s_{3}$ x9 1.00 0.61 0.74 0.15 1.00 0.66 1.00 0.89